Two charges separated by 1 m exert a 1.0-N force on each other. If the charges are pushed to 0..50-m separation, what will be the force on each charge?
1
Expert's answer
2018-09-10T12:42:08-0400
We can find the force on each charge from the Coulomb’s Law:
F=k (q_1 q_2)/r^2 ,
here, F is the force that act on each charge, q_1, q_2 is the two charges, k is the Coulomb’s constant, r is the separation between two charges.
Since, k, q_1 and q_2 are constants we can write:
F=k (q_1 q_2)/r^2 ~1/r^2 .
Then, if the charges are pushed to 0.5 m separation, the force on each charge will be:
F~1/r^2 =1/(0.5 m)^2 =4.0 N.